Tickets to a concert are sold according to a poisson distribution with mean 30 per day.What are the probability that

i) less than 20 tickets are sold in one day

ii)all 180 tickets are sold in 5 day working week


1
Expert's answer
2021-08-20T09:55:13-0400

Poisson distribution with "\lambda=30" per day.

"p(x)=\frac{e^{-\lambda} \lambda^x}{x!}, x=0,1,2,..."

i. P(x<20)

"P(x<20)=\sum_{x=0}^{19}\frac{e^{-30} 30^x}{x!}"

The table below gives the values for x = 0-19


From the table, "\sum_{x=0}^{19}=0.02187"

Thus, "P(x<20)=0.02187"

=POISSON.DIST(19,30,TRUE) excel function can be used to obtain the solution.

ii. P(x=180) in 5 days

"\lambda=30\times5=150"

"p(180)=\frac{e^{-150} 150^{180}}{180!}"

"=0.00178"